OpenStudy (anonymous):

I am trying to figure out how to solve this proportion, 4x^2/ 8x^2 - 8x = 2/x, I am stuck, I got an answer, but I checked it and it was wrong. Does anyone know how to do it? If so, can u please show me the work?

OpenStudy (anonymous):

Solving for x

OpenStudy (phi):

\[ \frac{4x^2}{8x^2 -8x} =\frac{2}{x} \] can you factor 8x from the denominator of the fraction on the left side ?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

Don't u first have to simplify the proportion?

OpenStudy (phi):

what do you get when you factor 8x from the denominator ?

OpenStudy (phi):

this is the first step in simplifying the left side

OpenStudy (anonymous):

I got x/-16x

OpenStudy (phi):

factor out 8x from \[ 8x^2 -8x \] what do you get ?

OpenStudy (phi):

notice both terms can be evenly divided by 8x

OpenStudy (anonymous):

0?

OpenStudy (phi):

do you know the distributive property: 8x( x -1) what do you get if you multiply the 8x times each term inside the parens ?

OpenStudy (anonymous):

8x^2 + 8x?

OpenStudy (phi):

yes. except it is 8x^2 -8x (with a minus sign) notice that 8x^2 - 8x can be written 8*x*x - 8*x notice that there is an 8*x in both terms, so you can "undo" the distributive property you get 8*x * ( x - 1) you now have \[ \frac{4x^2}{8x(x-1)} = \frac{2}{x} \] notice that you can simplify the left. what is 4/8 ?

OpenStudy (anonymous):

1/2

OpenStudy (phi):

what is x^2/x or x*x/x ?

OpenStudy (anonymous):

X

OpenStudy (phi):

that means you now have \\[ \frac{\cancel{4x^2}x}{\cancel{8x}2(x-1)} = \frac{2}{x} \\ \frac{x}{2(x-1)} =\frac{2}{x}\] do you know how to cross multiply?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

\frac{\cancel{4x^2}x}{\cancel{8x}2(x-1)} = \frac{2}{x} \\ \frac{x}{2(x-1)} =\frac{2}{x}\], I don't get what this means

OpenStudy (phi):

I was trying to show that the 4x^2/8x simplifies to 1x/2 or just x/2

OpenStudy (anonymous):

I got 4 = x^2

OpenStudy (anonymous):

X^2 - 4

OpenStudy (phi):

if you do it step by step \[ \frac{4x^2}{8x} = \frac{1 x^2}{2x} \] using 4/8= 1/2 and then \[ \frac{1 x^2}{2x} = \frac{1 x}{2} = \frac{x}{2}\] using x^2/x = x

OpenStudy (anonymous):

X = 2, x = -2?

OpenStudy (phi):

can you show how you did the cross multiply?

OpenStudy (anonymous):

i cross multiplied x/2 and 2/x and got x^2 = 4

OpenStudy (phi):

that is correct, except that your problem is \[ \frac{x}{2(x-1)} =\frac{2}{x} \]

OpenStudy (anonymous):

I got 8x + 8 = x^2

OpenStudy (phi):

show your steps

OpenStudy (anonymous):

I multiplied 2x + 2 by 2 and get 4x + 4?

OpenStudy (phi):

ok, but there is no 2x+2 Do you mean 2(x-1) = 2x - 2 ?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

4x -4

OpenStudy (phi):

what do you get when you cross multiply \[ \frac{x}{2(x-1)} =\frac{2}{x} \]

OpenStudy (anonymous):

4x - 4 = x^ 2

OpenStudy (phi):

ok, now move all the terms to the x^2 side. add -4x to both sides add +4 to both sides what do you get ?

OpenStudy (anonymous):

So, x = 2?

OpenStudy (phi):

are you a human calculator? we have to solve a quadratic equation here. 4x - 4 = x^ 2 the first step is put it in standard form. add -4x to both sides add +4 to both sides what do you get ?

OpenStudy (anonymous):

x^2 - 4x + 4

OpenStudy (phi):

what happened to the = sign and the other side ?

OpenStudy (phi):

you should get x^2 - 4x + 4 = 0 now you have to factor this. Can you do that ?

OpenStudy (anonymous):

Yes, I got x = 2 and x = 2

OpenStudy (phi):

to show your work: x^2 - 4x + 4 = 0 (x-2)(x-2) =0 so either (x-2)=0 or (x-2)=0 (we get "repeated" roots) x-2=0 x= 2 (twice)

OpenStudy (anonymous):

Okay, thanks a bunch!